# A Concentration Bound for TD(0) with Function Approximation

### v3

@@ Abstract
We derive a concentration bound of the type ‘for all $n\geq n_{0}$ for some $n_{0}$’ for TD(0) with linear function approximation. We work with online TD learning with samples from a single sample path of the underlying Markov chain. This makes our analysis significantly different from offline TD learning or TD learning with access to independent samples from the stationary distribution of the Markov chain. We treat TD(0) as a contractive stochastic approximation algorithm, with both martingale and Markov noises. Markov noise is handled using the Poisson equation and the lack of almost sure guarantees on boundedness of iterates is handled using the concept of relaxed concentration inequalities.
@@ 1 Introduction
TD(0) is one of the most popular reinforcement learning algorithms for policy evaluation (Tsitsiklis and Van Roy 1997). Given a fixed policy, the algorithm is an iterative method to obtain the value function for each state under the long-term discounted reward framework. To mitigate the issues of large state spaces, the value function is often approximated using a linear combination of feature vectors. This algorithm is referred to as TD(0) with linear function approximation. In this paper, we work with online TD(0) with a single sample path of the underlying Markov chain. Our goal in this paper is to obtain a concentration bound of the form from some time on or more precisely, for all $n\geq n_{0}$ for a suitably chosen $n_{0}$ for this algorithm.
@@
A bound of this form was published for TD(0) as a section in our paper titled “Concentration of Contractive Stochastic Approximation and Reinforcement Learning” (Chandak et al. 2022). This paper gave an all time bound for contractive stochastic approximation with Markov noise and applied the bound to asynchronous Q-learning and TD(0). Although the main theorem and its application to asynchronous Q-learning are correct, TD(0) does not satisfy a key assumption for the main theorem [^3] In Chandak et al. (2022), TD(0) does not satisfy assumption (6) which is required in the proof for Lemma 1 that shows almost sure boundedness of the iterates. The authors thank Zaiwei Chen for pointing this out. and hence the theorem was incorrectly applied to TD(0). We remove the need for that assumption in this version, giving a completely different proof specific to the TD(0) algorithm.
@@
The previous paper required the iterates of the stochastic approximation iteration to be bounded by a constant with probability 1. This assumption is not known to be true for the iterates of online TD(0) with function approximation for a single sample path. In fact, a common method to alleviate this issue is to project the iterates back into a ball centered around the origin (Bhandari et al. 2018, Patil et al. 2023). The key difficulty caused by the lack of this assumption is in applying martingale inequalities which often require some restrictions on the increments of the martingale that are often not easy to verify. We do not work with the projection operator, and instead adapt relaxed concentration inequalities (Chung and Lu 2006, Section 8) for our problem. These bounds have an extra term given by the probability of increments going above a certain threshold (Tao and Vu 2015, Prop. 34).
@@
While there is a considerable body of work on concentration inequalities for reinforcement learning algorithms, these always aim for a moment bound or a high probability bound for sufficiently large time $n$. In contrast, we look at the complementary object, viz., concentration for all time after some $n_{0}\geq 0$. Moment bounds for TD(0) have been established, e.g., in Srikant and Ying (2019), Bhandari et al. (2018), and Chen et al. (2021). Other types of bounds (e.g., finite sample bounds with high probability results) have been established for multiple different settings in Prashanth et al. (2021) (consider uniform sampling from dataset), Patil et al. (2023) (consider projection and tail averaging), and Chen et al. (2025), Dalal et al. (2018a) (consider i.i.d. samples of state - action - next state triplets $(s,a,s^{\prime})$).
Going beyond TD(0), Bhandari et al. (2018), Chen et al. (2020, 2021), Wang et al. (2020), Dalal et al. (2018a, b), Li et al. (2020, 2023), Even-Dar and Mansour (2004), Prashanth et al. (2021), Qu and Wierman (2020), Sidford et al. (2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for various classes of algorithms. There is also a parallel body of work which seeks bound, either finite time or asymptotic, on the difference between the value function under the learned policy and the optimal value function (Jin et al. 2018, Yang et al. 2020, Yang and Wang 2019), i.e., the ‘regret’. Once again, this is distinct from our objective, which is to obtain a high probability bound valid for all time from some time on.
@@
Inspired by Chen et al. (2020), one of us considered stochastic approximation involving contractive maps and martingale noise, and derived such concentration bounds ‘from some $n_{0}$ on’ for this class of algorithms (Borkar 2021). This covered in particular synchronous Q-learning for discounted cost and some related schemes. In Chandak et al. (2022), we extended this work to cover ‘Markov noise’ (Meerkov 1972) in the stochastic approximation scheme, allowing us to give bounds for the asynchronous case. As mentioned before, this work assumed almost-sure boundedness of iterates, which is not satisfied by the TD(0) algorithm. We remove the need for this assumption in our current work. Other articles aiming at bounds of these forms, such as the one we provide, are found in Chandak et al. (2023) for the LSPE algorithm, and Borkar (2002), Thoppe and Borkar (2019), Kamal (2010) for abstract stochastic approximation schemes. Another work which considers all-time bounds is Chen et al. (2025), but they only consider additive and multiplicative noise and not Markovian noise as considered in this paper.
To summarize, conventional concentration or sample complexity bounds yield bounds, e.g. mean square or ‘with high probability’ bounds, on the error from the target after $n$ iterations starting from time zero. The asymptotic regret bounds get an upper bound or lower bound or both on how some measure of cumulative error grows with time in an asymptotic sense. Our bounds differ from both. They are ‘all time bounds’ in the sense that they give a high probability bound for the iterates to remain in a prescribed small neighborhood of the target for all time from some time $n_{0}$ onwards. In fact the requirement ‘from some $n_{0}$ onwards’ is dictated by the fact that the decreasing stepsize needs to be sufficiently small from $n_{0}$ on. As for regret bounds, these are for cumulative error and are typically ‘almost sure’ or ‘in mean’ and asymptotic in nature unlike our bounds, which are from some time on, but with high probability.
@@
The rest of the paper is structured as follows. Section 2 gives a background to the TD(0) algorithm, and states the main result. The result is proved in Section 3. A concluding section highlights some future directions. Appendix A states a martingale inequality used in our proof and Appendix B gives proofs for some technical lemmas which are used to prove the main theorem.
1.1 Notation
@@
2 TD(0) - Background and Main Result
TD(0) is an algorithm for policy evaluation, i.e., for learning the performance of a fixed policy, and not for optimizing over policies (Tsitsiklis and Van Roy 1997). Hence a stationary policy is fixed a priori, giving us a time homogeneous uncontrolled Markov chain $\{Y_{n}\}$ over a finite state space $S$. The transition probabilities are given by $p(\cdot|\cdot)$, where the dependence on the policy is suppressed. Assume that the chain is aperiodic irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(s)],s=|S|$. Let $D$ denote the $s\times s$ diagonal matrix whose $i$th diagonal entry is $\pi(i)$. Reward $r(i)$ is received when a transition from state $i$ takes place. The objective is to evaluate the long term discounted reward for each state given by the “value function”
$V(i)=E\left[\sum_{m=0}^{\infty}\gamma^{m}r(X_{m})|X_{0}=i\right],i\in S.$
Here $10<\gamma<$ 1 is a discount factor. The dynamic programming equation for evaluating the same is given by
$V(i)=r(i)+\gamma\sum_{j}p(j|i)V(j),\ i\in S,$
which can be written as the following vector equation
$V=r+\gamma PV$
for $r=[r(1),...,r(s)]^{T}$ and $P=[[p(j|i)]]_{i,j\in S}\in\mathcal{R}^{s\times s}$.
The state space can often be large ($s\gg 1$), and to alleviate this ‘curse of dimensionality’, $V$ is often approximated using a linear combination of $d$ linearly independent basis functions (feature vectors) $\phi_{i}:S\mapsto\mathcal{R},1\leq i\leq d$, with $s>>d\geq 1$. Also, let $\varphi(i)=[\phi_{1}(i),\ldots,\phi_{d}(i)]^{T}$. Thus $V(i)\approx\sum_{j=1}^{d}x(j)\phi_{j}(i)$, i.e., $V(i)=x^{T}\varphi(i)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $s\times d$ matrix whose $i$th column is $\phi_{i}$. Here $x$ denotes the learnable weights for the linear function approximator. Since $\{\phi_{i}\}$ are linearly independent, $\Phi$ is full rank. Substituting this approximation into the dynamic programming equation above leads to
@@ 2 TD(0) - Background and Main Result
where $\Pi$ denotes the projection to Range($\Phi$) with respect to a suitable norm. It turns out to be convenient to take projection with respect to the weighted norm $\|x\|_{D}\coloneqq\sqrt{x^{T}Dx}=(\sum_{i}\pi(i)|x(i)|^{2})^{1/2}$ whereby the projection map is
$\Pi x\coloneqq\Phi(\Phi^{T}D\Phi)^{-1}\Phi^{T}Dx.$
@@
Recall that $\varphi(i)^{T}\in\mathcal{R}^{d}$ for $i\in S$ denotes the $i$th row of $\Phi$ and $a(n)$ denotes the positive stepsize sequence. We now make the following two assumptions:
@@
    First we make an assumption on $\Phi$. For that, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the square of the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$. Assume that
@@
    Since the feature vectors can be scaled without affecting the algorithm (the weights $x(i)$ get scaled accordingly), this assumption does not restrict the algorithm. Alternatively, the stepsize can be appropriately scaled as well, i.e., $a(n)=b(n)c$, where $b(n)$ acts as the effective stepsize and $c\varphi(\cdot)$ act as the effective feature vectors which satisfy the above assumption. This assumption can be replaced with the following assumption on the basis vectors.
    $\|\varphi(i)\|\leq\frac{\sqrt{2(1-\gamma)}}{1+\gamma}\;\forall i,$
    i.e., the $\ell_{2}$ norm of each row of $\Phi$ is bounded by $\sqrt{2(1-\gamma)}/1+\gamma$. To see this, note that
    $\frac{\|\Psi x\|_{2}}{\|x\|_{2}}=\frac{\|\Phi x\|_{D}}{\|x\|_{2}}\leq\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}.$
    Now $\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}$ is the operator norm defined with $\ell_{2}$ norm for domain and $\ell_{\infty}$ for co-domain, which is equal to the maximum $\ell_{2}$ norm of a row. Hence
    $\lambda_{M}=\max_{x\neq\theta}\frac{\|\Psi x\|_{2}}{\|x\|_{2}}\leq\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}=\max_{i}\|\varphi(i)\|_{2}.$
@@
We next rearrange iteration (2) to separate the martingale noise and the ‘Markov noise’.
@@
$=$ $x_{n}+a(n)\left(\sum_{i\in S}\pi(i)F(x_{n},i)-x_{n}\right)+\underbrace{a(n)M_{n+1}x_{n}}_{\tau_{1}}+\underbrace{a(n)\left(F(x_{n},Y_{n})-\sum_{i\in S}\pi(i)F(x_{n},i)\right)}_{\tau_{2}}.$
@@
$F(x,Y)=\varphi(Y)r(Y)+\gamma\varphi(Y)\sum_{j}p(j|Y)\varphi(j)^{T}x-\varphi(Y)\varphi(Y)^{T}x+x$
@@
$M_{n+1}=\gamma\varphi(Y_{n})\Big{(}\varphi(Y_{n+1})^{T}-\sum_{j}p(j|Y_{n})\varphi(j)^{T}\Big{)}.$
@@
where $\theta$ denotes the zero vector. Then term $\tau_{1}$ denotes the error due to the martingale noise term and term $\tau_{2}$ denotes the error due to the ‘Markov noise’ $\{Y_{n}\}$.
The following lemma shows that the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction. Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$. Note that the following lemma shows that $\sum_{i}\pi(i)F(\cdot,i)$ is contractive under the Euclidean norm with an explicit expression for the contraction factor.
@@
Moreover, $10<\alpha<$ 1 and hence the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction.
The proof appears in Appendix B. The Banach contraction mapping theorem implies that $\sum_{i}\pi(i)F(\cdot,i)$ has a unique fixed point $x^{*}$, i.e., there exists a unique point $x^{*}$ such that $\sum_{i}\pi(i)F(x^{*},i)=x^{*}$. Then,
$\sum_{i}\pi(i)F(x^{*},i)=(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)x^{*}=x^{*}$
@@
For $n\geq 0$, we further define:
[…]

### v4

@@ Abstract
We derive uniform all-time concentration bound of the type ‘for all $n\geq n_{0}$ for some $n_{0}$’ for TD(0) with linear function approximation. We work with online TD learning with samples from a single sample path of the underlying Markov chain. This makes our analysis significantly different from offline TD learning or TD learning with access to independent samples from the stationary distribution of the Markov chain. We treat TD(0) as a contractive stochastic approximation algorithm, with both martingale and Markov noises. Markov noise is handled using the Poisson equation and the lack of almost sure guarantees on boundedness of iterates is handled using the concept of relaxed concentration inequalities.
@@ 1 Introduction
TD(0) is one of the most popular Reinforcement Learning (RL) algorithms for policy evaluation (Tsitsiklis and Van Roy 1997). Given a fixed policy, the algorithm is an iterative method to obtain the value function for each state under the long-term discounted reward framework. To mitigate the issues of large state spaces, the value function is often approximated using a linear combination of feature vectors. This algorithm is referred to as TD(0) with linear function approximation. In this paper, we work with online TD(0) with a single sample path of the underlying Markov chain. Our goal in this paper is to obtain a concentration bound of the form from some time on or more precisely, for all $n\geq n_{0}$ for a suitably chosen $n_{0}$ for this algorithm.
@@
A bound of this form was published for TD(0) as a section in our paper titled “Concentration of Contractive Stochastic Approximation and Reinforcement Learning” (Chandak et al. 2022). This paper established an all-time bound for contractive stochastic approximation with Markov noise and applied the bound to asynchronous Q-learning and TD(0). Although the main theorem and its application to asynchronous Q-learning are correct, TD(0) does not satisfy a key assumption for the main theorem [^3] In Chandak et al. (2022), TD(0) does not satisfy assumption (6) which is required in the proof for Lemma 1 that shows almost sure boundedness of the iterates. The authors thank Zaiwei Chen for pointing this out. and hence the theorem was incorrectly applied to TD(0). We remove the need for that assumption in this version, giving a completely different proof tailored to the TD(0) algorithm.
@@
The previous paper required the iterates of the stochastic approximation iteration to be bounded by a constant with probability 1. This assumption is not known to be true for the iterates of online TD(0) with function approximation for a single sample path. In fact, a common method to alleviate this issue is to project the iterates back into a ball centered around the origin (Bhandari et al. 2018, Patil et al. 2023). The key difficulty caused by the lack of this assumption is in applying martingale inequalities, which often require some restrictions on the increments of the martingale that are often not easy to verify. We do not modify the algorithm and instead adapt relaxed concentration inequalities (Chung and Lu 2006, Section 8) for our problem. These bounds have an extra term given by the probability of increments going above a certain threshold (Tao and Vu 2015, Prop. 34). While the proof in this paper is restricted to TD(0), the underlying idea of using relaxed concentration inequalities is broadly applicable to other algorithms that face similar challenges due to unboundedness.
1.1 Related Works
There has been growing interest in analyzing the finite-time performance of reinforcement learning (RL) algorithms. Existing results can broadly be categorized by the type of bounds they establish. The most extensive body of work concerns expectation or mean-square bounds (see, e.g., Chen et al. (2020, 2021)). Another prominent line of research focuses on regret bounds, which characterize how the cumulative error grows over time—typically through almost-sure or expected regret guarantees (see, e.g., Azar et al. (2017), Jin et al. (2018), Yang et al. (2020), Yang and Wang (2019)). A third class comprises high-probability or concentration bounds (see, e.g., Li et al. (2023), Qu and Wierman (2020), Even-Dar and Mansour (2004). Our work falls within this category but differs from conventional analyses that establish high-probability guarantees only for sufficiently large time $n$. In contrast, we derive uniform all-time bounds, i.e., bounds that hold for all $n\geq n_{0}$ with probability at least $1-\delta$.
@@ 1.1 Related Works
Specifically for TD(0), moment bounds have been established in Srikant and Ying (2019), Bhandari et al. (2018), and Chen et al. (2021). High probability bounds have been established under various modifications of the TD(0) algorithm. These include uniform sampling from dataset (Prashanth et al. 2021), projection and tail averaging (Patil et al. 2023), and oracle access to i.i.d. samples of state - action - next state triplets $(s,a,s^{\prime})$ (Chen et al. 2025, Dalal et al. 2018).
@@
One of us considered stochastic approximation involving contractive maps and martingale noise, and derived maximal concentration bounds for this class of algorithms (Borkar 2021). This covered, in particular, synchronous Q-learning for discounted cost and some related schemes. In Chandak et al. (2022), we extended this work to cover ‘Markov noise’ (Meerkov 1972) in the stochastic approximation scheme, allowing us to give bounds for the asynchronous case. As mentioned before, this work assumed almost-sure boundedness of iterates, which is not satisfied by the TD(0) algorithm. We remove the need for this assumption in our current work. Other articles aiming at bounds of these forms, such as the one we provide, are found in Chandak et al. (2023) for the LSPE algorithm, and Borkar (2002), Thoppe and Borkar (2019), Kamal (2010) for abstract stochastic approximation schemes. A recent work (Chen et al. 2025) considers all-time bounds for iterates without almost-sure boundedness, but they only consider additive and multiplicative noise and not Markovian noise as considered in this paper. Their proof technique relies on Moreau envelopes and a bootstrapping technique, which differs significantly from our work.
1.2 Outline and Notation
@@
The rest of the paper is structured as follows. Section 2 gives a background to the TD(0) algorithm, along with the required assumptions and the stochastic approximation formulation. Section 3 states the main result and provides some insights into the result. The result is proved in Section 4. A concluding section highlights some future directions. Appendix A states a martingale inequality used in our proof and Appendix B gives proofs for some technical lemmas which are used to prove the main theorem.
@@
2 Background on TD(0)
TD(0) is an algorithm for policy evaluation, i.e., for learning the performance of a fixed policy, and not for optimizing over policies. Hence a stationary policy is fixed a priori, giving us a time homogeneous uncontrolled Markov chain $\{Y_{n}\}$ over a finite state space $\mathcal{S}$. The transition probabilities are given by $p(\cdot|\cdot)$, where the dependence on the policy is suppressed. Assume that the chain is aperiodic irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(S)],S=|\mathcal{S}|$. Let $D$ denote the $S\times S$ diagonal matrix whose $s$th diagonal entry is $\pi(s)$. Reward $r(s)$ is received when a transition from state $s$ takes place. Note that this reward can be stochastic as well, and the additional noise thereof can be combined with other noise terms without affecting our concentration result. For simplicity, we assume that we receive a deterministic $r(s)$. The objective is to evaluate the long term discounted reward for each state given by the “value function”
$V(s)=E\left[\sum_{m=0}^{\infty}\gamma^{m}r(X_{m})|X_{0}=s\right],s\in\mathcal{S}.$
Here $0<\gamma<1$ is the discount factor. The dynamic programming equation for evaluating the same is
$V(s)=r(s)+\gamma\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)V(s^{\prime}),\ s\in\mathcal{S}.$
This can be written as the following vector equation
$V=r+\gamma PV,$
for $r=[r(1),...,r(S)]^{T}$ and $P=[[p(s^{\prime}|s)]]_{s^{\prime},s\in\mathcal{S}}\in\mathbb{R}^{S\times S}$.
The state space can often be large ($S\gg 1$), and to alleviate this ‘curse of dimensionality’, $V$ is often approximated using a linear combination of $d$ linearly independent basis functions (feature vectors) $\phi_{i}\in\mathbb{R}^{S},1\leq i\leq d$, with $S>>d\geq 1$. Also, let $\varphi(s)=[\phi_{1}(s),\ldots,\phi_{d}(s)]^{T}$ for $s\in\mathcal{S}$ denote the vector comprising of components corresponding to state $s$ in each feature. Thus $V(s)\approx\sum_{i=1}^{d}x(i)\phi_{i}(s)$, i.e., $V(s)\approx x^{T}\varphi(s)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $S\times d$ matrix whose $i$th column is $\phi_{i}$. Here $x$ denotes the learnable weights for the linear function approximator. Since $\{\phi_{i}\}$ are linearly independent, $\Phi$ is full rank. Substituting this approximation into the dynamic programming equation above leads to
@@ 2 Background on TD(0)
where $\Pi$ denotes the projection to Range($\Phi$) with respect to a suitable norm. It turns out to be convenient to take projection with respect to the weighted norm $\|y\|_{D}\coloneqq\sqrt{y^{T}Dy}=(\sum_{s\in\mathcal{S}}\pi(s)(y(s))^{2})^{1/2}$ for $y\in\mathbb{R}^{S}$. The projection map with respect to this norm is
$\Pi y\coloneqq\Phi(\Phi^{T}D\Phi)^{-1}\Phi^{T}Dy.$
@@
Here $a(n)$ denotes the positive stepsize sequence. At the end of this section, we explain how this iteration can be expected to converge to the required fixed point from (1).
2.1 Assumptions
We impose two assumptions on the algorithm. The first is about the feature vectors, and as we explain next, it does not restrict the algorithm. The second specifies the class of stepsizes $a(n)$ considered, which is standard in analysis of RL. In fact, our results hold for a broader class of step sizes than those typically required in stochastic approximation frameworks.
@@ 2.1 Assumptions
    For the assumption on $\Phi$, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the square of the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$. Assume that
@@
    Since the feature vectors can be scaled without affecting the algorithm (the weights $x(i)$ get scaled accordingly), this assumption does not restrict the algorithm. Alternatively, the stepsize can be appropriately scaled as well, i.e., $a(n)=b(n)c$, where $b(n)$ acts as the effective stepsize and $c\varphi(\cdot)$ act as the effective feature vectors which satisfy the above assumption. This assumption can be replaced with the following assumption on the basis vectors. ∥φ(s)∥≤2(1-γ)1+γ ∀s∈S, i.e., the $\ell_{2}$ norm of each row of $\Phi$ is bounded by $\sqrt{2(1-\gamma)}/(1+\gamma)$. To see this, note that ∥Ψx∥2∥x∥2=∥Φx∥D∥x∥2≤∥Φx∥∥x∥2.Now $\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}$ is the operator norm defined with $\ell_{2}$ norm for domain and $\ell_{\infty}$ for co-domain, which is equal to the maximum $\ell_{2}$ norm of a row. Hence λ_M=max_x≠θ∥Ψx∥2∥x∥2≤max_x≠θ∥Φx∥∥x∥2=max_s∈Sφ(s)∥_2.
@@
2.2 Formulation as a Stochastic Approximation Iteration
We next rearrange algorithm (2) to separate the martingale noise and the ‘Markov noise’, and write it as a stochastic approximation iteration.
@@ 2.2 Formulation as a Stochastic Approximation Iteration
$=$ $x_{n}+a(n)\left(\sum_{s\in\mathcal{S}}\pi(s)F(x_{n},s)-x_{n}\right)+\underbrace{a(n)M_{n+1}x_{n}}_{\tau_{1}}+\underbrace{a(n)\left(F(x_{n},Y_{n})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{n},s)\right)}_{\tau_{2}},$
@@
$F(x,Y)=\varphi(Y)r(Y)+\gamma\varphi(Y)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y)\varphi(s^{\prime})^{T}x-\varphi(Y)\varphi(Y)^{T}x+x,$
@@
$M_{n+1}=\gamma\varphi(Y_{n})\Big(\varphi(Y_{n+1})^{T}-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{n})\varphi(s^{\prime})^{T}\Big).$
@@
where $\theta$ denotes the zero vector. The term $\tau_{1}$ denotes the error due to the martingale noise term and term $\tau_{2}$ denotes the error due to the ‘Markov noise’ $\{Y_{n}\}$.
The following lemma shows that the function $\sum_{s\in\mathcal{S}}\pi(s)F(\cdot,s)$ is a contraction. Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$. While the contraction property of the TD(0) algorithm is well known Tsitsiklis and Van Roy (1997), we obtain an explicit expression for the contraction factor.
@@
Moreover, $0<\alpha<1$ and hence the function $\sum_{s\in\mathcal{S}}\pi(s)F(\cdot,s)$ is a contraction.
The proof appears in Appendix B. The Banach contraction mapping theorem implies that $\sum_{s\in\mathcal{S}}\pi(s)F(\cdot,s)$ has a unique fixed point $x^{*}$, i.e., there exists a unique point $x^{*}$ such that $\sum_{s\in\mathcal{S}}\pi(s)F(x^{*},s)=x^{*}$. We next show that the fixed point $x^{*}$ is the required fixed point we wish to converge to. Before that, we first observe that
$\sum_{s\in\mathcal{S}}\pi(s)F(x,s)=(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)x.$
Then,
$\sum_{s\in\mathcal{S}}\pi(s)F(x^{*},s)=(\Phi^{T}Dr+\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi+I)x^{*}=x^{*}$
@@
3 Main Result
Before stating the main result, we define the following two sequences. For $n\geq 0$,
$b_{k}(n)$ $=$ $\sum_{m=k}^{n}a(m),\ 0\leq k\leq n<\infty,$